OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y =____x +____

OpenStudy (eric_d):

Use y-y1=m(x-x1)

OpenStudy (eric_d):

Then, rearrange it into y=mX+c

OpenStudy (eric_d):

@Briinndyy_69

OpenStudy (anonymous):

Im still confused.. im really bad at math